# A Calculus of Overlays

### v2

@@ 8. Future Work
The Overlay language (Yang, 2025b) is an executable implementation of overlay-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the inheritance, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language, spanning both theoretical foundations and library-level encodings.
Type system
The Overlay language’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how overlay-calculus relates to the $\lambda$-calculus: a Typed overlay-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with overlay inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that a inheritance provides implementations for all required slots, not merely that references resolve. In untyped overlay-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
Standard library
@@
Overlay-calculus is open by default: inheritance can freely merge any two overlays, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is not a defect; it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5). However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one constructor. The observer pattern (Section 5) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
@@ 9. Conclusion
The goals of declarative programming forced the semantics of overlay-calculus upon us. The emergent phenomena exhibit an expressive asymmetry between overlay-calculus and the $\lambda$-calculus. Inheritance is not a mechanism for organizing functions; functions are syntactic sugar for inheritance.
@@ Appendix A Expressive Asymmetry: Proofs
The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into overlay-calculus.
@@
Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an overlay-calculus expression that depends only on the construct’s immediate subexpressions. No rule inspects the surrounding context. ∎
Theorem A.3 (Non-expressibility of overlay).
There is no macro-expressible translation from overlay-calculus to the $\lambda$-calculus.
@@
Let $\mathscr{L}_{1}$ be the overlay-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of overlay trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$. Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$. Define $\cong_{1}$ on $\mathscr{L}_{1}$ as overlay-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every overlay context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
@@
In overlay-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$. Define overlay trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
@@
$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
@@
In the composite scope $\{S,\,C\}$, the definition of $\mathrm{eq}$ gains $\mathrm{repr}$ as a new sibling of $x_{1},x_{2},x_{3}$ inside the $\lambda b$ scope. By open recursion, the de Bruijn references $\uparrow^{\!2}.\mathrm{argument}$ and $\uparrow^{\!1}.\mathrm{argument}$ resolve through the same scope chain as the existing let-bindings. Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$. These produce structurally distinct overlay trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$. By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the overlay facilities of $\mathscr{L}_{1}$. ∎
@@
Overlay-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
@@ Appendix B Multi-Path Self-Reference in Other Systems
This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that overlay-calculus handles naturally.
@@
object MyObjectA extends Object1.MyInner
@@
Scala 3 rejects MyObjectA with the error:
@@
object MyObjectA cannot be instantiated since
@@
The equivalent overlay-calculus definition is:
@@
$\mathrm{MyObjectA}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
This is well-defined in overlay-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{MyObjectA}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{MyObjectA})$ for overlay paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{MyObjectA}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties and no disambiguation is needed.
@@
        # object MyObjectA extends
@@
        options.MyObjectA = lib.mkOption {
@@
  builtins.attrNames result.config.MyObjectA.outer
@@
error: The option ‘MyObjectA.outer’
@@
The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In overlay-calculus, the same pattern is well-defined because $\mathtt{overlays}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
@@ C.1. The Product Lattice
$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–overlay pairs)
$D_{\mathtt{overlays}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
@@
Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$. The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overlays}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
@@ C.2. The Operator
Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overlays},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation. The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$. If $\mathtt{bases}\sqsubseteq\mathtt{bases}^{\prime}$ (pointwise), then $\mathtt{bases}^{*}\sqsubseteq\mathtt{bases}^{\prime*}$ (pointwise), since enlarging the one-step relation can only enlarge the transitive closure.
@@ C.3. Monotonicity
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\,\bigr\}.$
@@
$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{overlay}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\,\bigr\}$. Both $\mathtt{bases}^{*}$ and $\mathtt{overlays}$ appear as sets being iterated over. As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$. Enlarging either $\mathtt{bases}$ or $\mathtt{overlays}$ can only add elements to the result. Positive in $\mathtt{bases}$ and $\mathtt{overlays}$.
Equation (3): $\mathtt{overlays}$
@@
$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
@@
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
Both $\mathtt{overlays}$ and $\mathtt{resolve}$ appear as sets being iterated over. The function $\mathtt{inherits}$ is an AST primitive. Enlarging either $\mathtt{overlays}$ or $\mathtt{resolve}$ can only add targets to the result. Positive in $\mathtt{overlays}$ and $\mathtt{resolve}$.
@@
For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $p_{\mathrm{overlay}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function. Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$. Positive in $\mathtt{supers}$.

### v4

@@ 8. Future Work
MIXINv2 (Yang, 2026) is an executable implementation of inheritance-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the mixin tree, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language, spanning both theoretical foundations and library-level encodings.
8.1. Type System
MIXINv2’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how inheritance-calculus relates to the $\lambda$-calculus: a Typed inheritance-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with mixin inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that an inheritance provides implementations for all required slots, not merely that references resolve. In untyped inheritance-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
8.2. Standard Library
@@
Inheritance-calculus is open by default: inheritance can freely merge any two mixins, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is not a defect; it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5). However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one constructor. The observer pattern (Section 5) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
@@ 9. Conclusion
Immunity to nonextensibility guided the design of inheritance-calculus. The emergent phenomena exhibit an expressive asymmetry between inheritance-calculus and the $\lambda$-calculus. Inheritance is not a mechanism for organizing functions; functions are syntactic sugar for inheritance.
@@ Appendix A Expressive Asymmetry: Proofs
The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into inheritance-calculus.
@@
Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an inheritance-calculus expression that depends only on the construct’s immediate subexpressions. No rule inspects the surrounding context. ∎
Theorem A.3 (Non-expressibility of inheritance-calculus).
There is no macro-expressible translation from inheritance-calculus to the $\lambda$-calculus.
@@
Let $\mathscr{L}_{1}$ be the inheritance-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of mixin trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$. Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$. Define $\cong_{1}$ on $\mathscr{L}_{1}$ as inheritance-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every inheritance-calculus context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
@@
In inheritance-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$. Define mixin trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
@@
$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{$
$\qquad\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
@@
In the composite scope $\{S,\,C\}$, the definition of $\mathrm{eq}$ gains $\mathrm{repr}$ as a new sibling of $x_{1},x_{2},x_{3}$ inside the $\lambda b$ scope. By open recursion, the de Bruijn references $\uparrow^{\!2}.\mathrm{argument}$ and $\uparrow^{\!1}.\mathrm{argument}$ resolve through the same scope chain as the existing let-bindings. Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$. These produce structurally distinct mixin trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$. By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the inheritance facilities of $\mathscr{L}_{1}$. ∎
@@
Inheritance-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
@@ Appendix B Multi-Path Self-Reference in Other Systems
This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that inheritance-calculus handles naturally.
@@
object HasMultipleOuters extends Object1.MyInner
@@
Scala 3 rejects HasMultipleOuters with the error:
@@
object HasMultipleOuters cannot be instantiated since
@@
The equivalent inheritance-calculus definition is:
@@
$\mathrm{HasMultipleOuters}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
This is well-defined in inheritance-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{HasMultipleOuters}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{HasMultipleOuters})$ for override paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties in this semantics, so no route-priority disambiguation rule is introduced.
@@
# object HasMultipleOuters extends
@@
options.HasMultipleOuters = lib.mkOption {
@@
builtins.attrNames result.config.HasMultipleOuters.outer
@@
error: The option HasMultipleOuters.outer
@@
The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In inheritance-calculus, the same pattern is well-defined because $\mathtt{overrides}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
@@ C.1. The Product Lattice
$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–override pairs)
$D_{\mathtt{overrides}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
@@
Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$. The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overrides}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
@@ C.2. The Operator
Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overrides},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation. The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$. If $\mathtt{bases}\sqsubseteq\mathtt{bases}^{\prime}$ (pointwise), then $\mathtt{bases}^{*}\sqsubseteq\mathtt{bases}^{\prime*}$ (pointwise), since enlarging the one-step relation can only enlarge the transitive closure.
@@ C.3. Monotonicity
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\,\bigr\}.$
@@
$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{override}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\,\bigr\}$. Both $\mathtt{bases}^{*}$ and $\mathtt{overrides}$ appear as sets being iterated over. As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$. Enlarging either $\mathtt{bases}$ or $\mathtt{overrides}$ can only add elements to the result. Positive in $\mathtt{bases}$ and $\mathtt{overrides}$.
Equation (3): $\mathtt{overrides}$
@@
$\mathtt{overrides}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
@@
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{override}}\in\mathtt{overrides}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{override}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
Both $\mathtt{overrides}$ and $\mathtt{resolve}$ appear as sets being iterated over. The function $\mathtt{inherits}$ is an AST primitive. Enlarging either $\mathtt{overrides}$ or $\mathtt{resolve}$ can only add targets to the result. Positive in $\mathtt{overrides}$ and $\mathtt{resolve}$.
@@
For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{override}}=p_{\mathrm{def}}\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $p_{\mathrm{override}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function. Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$. Positive in $\mathtt{supers}$.